# Towards Safe and Robust Autonomous Vehicle Platooning: A Self-Organizing Cooperative Control Framework

### v1

@@ III-A Problem Formulation and Reinforcement Learning Setup
    where $w_{C}$, $w_{L}$, $w_{F}$, and $w_{A}$ are the weights for the single-vehicle rewards: collision reward $r_{C}$, lane reward $r_{L}$, high-speed reward $r_{F}$, and acceleration reward $r_{A}$, respectively. Since collisions are unacceptable, the absolute value of $w_{C}$ will be significantly greater than the other coefficients. $w_{M}$, $w_{D}$, $w_{H}$, and $w_{S}$ correspond to the multi-vehicle collaborative rewards: same-lane reward $r_{M}$, distance traveled reward $r_{D}$, headway reward $r_{H}$, and speed coordination reward $r_{S}$. Single-vehicle rewards primarily ensure the safety and efficiency of individual vehicles, while multi-vehicle rewards maintain the overall structure of the platoon, preventing disbandment issues. We considere both individual vehicle rewards and multi-vehicle collaborative rewards. For individual vehicle rewards, the focus is on safety and efficiency. In contrast, multi-vehicle collaborative rewards emphasize the cooperation and coordination among vehicles. The schematic diagram of the reward function is shown in Figure 8.
    Figure 4: Illustration of the designed headway maintaining reward, which considers both the single-vehicle rewards and the multi-vehicle rewards
III-B Deep Reinforcement Learning Algorithm
Traditional DRL algorithms such as Deep Q-Networks (DQN) have been applied to autonomous driving decision-making[44]. However, these algorithms primarily focus on single-agent learning and perform inadequately in scenarios requiring complex multi-vehicle cooperative decision-making. This study addresses the challenge of vehicle platooning reconstruction in mixed traffic flow, aiming to develop a robust self-organizing decision model for vehicle platooning under highly disruptive conditions through agent collaboration.
To this end, we have adapted the PPO algorithm[45] to enhance its application to the safe and robust multi-vehicle cooperative decision-making problem. PPO is distinguished by its balance of simplicity, sample efficiency, and robustness, and has shown exceptional performance in various tasks.
In our approach, we utilize a CTDE strategy. As illustrated in Figure 5, the sensor data from all platooning vehicles are aggregated into a unified observation space. A shared policy $\pi_{\theta}$ then outputs the actions for all vehicles within the platoon. Following the execution of these actions by the low-level controllers, the agents update the policy $\pi_{\theta}$ based on rewards obtained from environmental feedback, employing experience replay and gradient-based optimization.
Figure 5: The illustration of the PPO-based formation configuration algorithm schematic model
@@ III-B Deep Reinforcement Learning Algorithm
In deep reinforcement learning, deep neural networks (DNNs) are used to approximate policies, enabling agents to learn how to make decisions within our environment to maximize cumulative rewards. DNNs play a crucial role in decision-making within deep reinforcement learning. This study employs the Actor-Critic architecture-based DRL neural network structure depicted in Figure 6.
Figure 6: Schematic diagram of the network structure based on the Actor-Critic architecture
@@
III-C Twin World Security Projection
While DRL enables robust decision-making in mixed traffic with strong interferences, its ”black-box” nature can lead to unexpected outcomes in standard scenarios and may not fully comply with traffic regulations. Although more reward functions can be used to impose these constraints, this approach can result in significant training costs and potential model divergence. This study addresses these issues by implementing a comprehensive and extensible safety mask.
To ensure decision safety, we propose a safety-priority-based twin-world model. This model evaluates the safety of DRL decisions by first detecting potential collisions based on safety priorities, thereby reducing computational complexity. It then recursively computes $T_{n}$ steps in a twin world constructed from the current scenario, optimizing actions based on the platoon’s safety margin.
Given the focus of this study on decision-making in mixed traffic, it is essential to predict the driving behavior of HDVs. To this end, we utilize the prevalent Intelligent Driver Model (IDM) [47] for simulating car-following behaviors. For lane-changing decisions, we adopt the MOBIL model [23]. These two models, when combined, ensure a minimal risk of collisions among HDVs while closely mirroring human driving patterns.
The safety mask models each vehicle as a circle with the vehicle’s center as the center and half the vehicle’s length as the radius. We simulate future $T_{n}$ steps using motion models of HDVs and CAVs to identify potential collision risks. While straightforward for single vehicles, this simulation becomes computationally challenging for vehicle platoons due to the exponentially increasing action space. To mitigate this, we propose a safety-priority-based simulation model. This model prioritizes vehicles with the smallest safety margins and conducts safety simulations in the twin world. The entire platoon tends to maintain cohesive formation, allowing for high computational efficiency suitable for real-time applications.
@@ 23      Replace action $a_{t}\leftarrow a_{t}^{\prime}$ Update $\phi[i]\leftarrow\phi[i+1],i=1,2,\cdots$
III-D Formation maintenance
The platoon reconfiguration model ensures safe and robust self-organizing decision-making and control in complex scenarios with strong interference. However, it struggles to maintain small headway distances in routine scenarios due to overly conservative strategies aimed at avoiding collision risks. To address this issue, we propose an adaptive switching mechanism, the platoon maintenance model, which integrates optimal control methods.
This model maintains stable, close headway distances in routine scenarios while dynamically adjusting driving strategies based on traffic conditions and interference levels. By monitoring the environment and vehicle states in real-time, the model adapts its strategies to ensure safety in complex scenarios and efficiency in routine conditions. This approach enhances overall coordination and driving efficiency, providing a flexible response to varying conditions.
To address the issue of decision control allocation in vehicle platooning across different scenarios, this study employs a Finite State Machine (FSM) for decision-making. At any given time, the state automaton can be in only one state. The FSM is defined by the five-tuple $M(Q,D,\delta,q_{0},F)$, where:
@@ III-D Formation maintenance
This study considers a scenario as a routine safe scenario when the platoon is on the same lane and there are no other traffic participants within $L_{safe}$ ahead or behind the platoon. In this case, the FSM switches to state $S_{1}$, adopting the LQR control model[48] to ensure smaller headway distances. In this study, string stability within the platoon[12] is not emphasized, as it has been thoroughly addressed in previous research. In all other scenarios, the FSM switches to state $S_{2}$, which employs a safety-enhanced data-driven strategy to balance the safety and efficiency of the platoon.
By defining these states and transitions, the FSM provides a structured approach to adaptively manage the control strategies of the platoon, ensuring optimal performance across varying traffic conditions. The detailed procedure is illustrated in Algorithm 2.
Input: $vehicles$, $riskLevel$, $Participants$, $L_{safe}$
Output: $controlStrategy$
1
2for $i=0$ to $N$ do
3       Assess risk level and number of traffic participants within $L_{safe}$;
5             Switch to State $S_{1}$: LQR Control Model;
6             Activate LQR Control;
7             Set $controlStrategy=LQR$;
8
9      else
10             Switch to State $S_{2}$: Safety-Enhanced Data-Driven Strategy;
11             Set $controlStrategy=DataDriven$;
12
13       end if
14      Execute $controlStrategy$ based on the current state;
15       if $controlStrategy==LQR$ then
16             ControlWithLQR();
17
18       end if
19      else if $controlStrategy==DataDriven$ then
20             ControlWithDataDrivenStrategy();
21
22       end if
23
24 end for
return $controlStrategy$
Algorithm 2 Adaptive Switching Mechanism
@@ IV-A1 Environment and simulation setting
The simulation environment is constructed upon the Highway-Env simulator[49], a reinforcement learning experimental environment library based on an OpenAI Gym environment[50]. As illustrated in Figure 3, all vehicles’ high-level policies are transformed into throttle and steering information through a low-level PID controller, thereby enabling vehicle movement within the simulation environment. Furthermore, our HDVs employ the IDM[47] car-following model and the MOBIL[23] lane-changing model provided by Highway-Env to achieve effective control in both longitudinal and lateral directions.
During the training process, a varying number of HDVs will randomly appear at randomly generated refresh points. The control sampling frequency is set to $15$ Hz, meaning decisions and controls are made 15 times per second.
We use three different random seeds and trained the proposed model in various complex scenarios, with each training session running for over 100,000 steps until the model converged. We selecte the Deep Q-Learning (DQN)[51] and Advantage Actor-Critic (A2C)[52] algorithms as experimental baselines. The training parameters for the DRL model are detailed in Table I.
@@
$w_{C}$ Weight for $r_{C}$200
$w_{F}$ Weight for $r_{F}$ 5
$w_{A}$ Weight for $r_{A}$ 1
$w_{M}$ Weight for $r_{M}$ 1.2
$w_{D}$ Weight for $r_{D}$ 1.2
$w_{H}$ Weight for $r_{H}$ 1
$w_{S}$ Weight for $r_{S}$ 2
@@ IV-A2 Performance evaluation
In this subsection, we evaluate the performance of the proposed model within a simulated training environment, comparing it with advanced RL algorithms (e.g., DQN, A2C) and traditional CACC and CLCC platooning strategies. Our results indicate that the proposed vehicle platooning model significantly outperforms existing models across various scenarios.
@@
We designed the scenario depicted in Figure 7 not merely as a training environment but as an edge case that stresses the safety and robustness of the model under extreme conditions. The platoon vehicles travel at a speed of 28 m/s within a mixed traffic flow, where surrounding vehicles’ speed distribution follows a uniform range of 25 to 28 m/s. The traffic flow is characterized by oscillations and diverse driving styles, including aggressive, conservative, and neutral behaviors, creating a volatile environment prone to accidents. Furthermore, we introduced a few malfunctioning vehicles that may suddenly stall or brake unexpectedly, further heightening the risk. The average inter-vehicle distance for HDVs is approximately 20 meters, providing insufficient space for the platoon to execute coordinated lane changes. Our study trained the model in this complex and perilous scenario, incorporating randomly generated traffic flows to simulate real-world conditions.
Figure 7: Schematic representation of the model training scenario wherein a multitude of interferences are present
  1. 1.
    Sample Efficiency: As shown in Figure 8, our safety-enhanced RL algorithm exhibits superior sample efficiency compared to baseline algorithms. The model achieves higher rewards per time step, rapidly increasing reward values within approximately 20,000 time steps and reaching initial convergence around 40,000 time steps. Despite typical training fluctuations, the model maintains an upward trend and stabilizes, indicating successful convergence. The consistently higher rewards across different scenarios are due to the model’s early prevention of illogical and illegal actions, reducing exploration time and enhancing operational efficiency.
    TABLE II: Testing performance comparison of the collision rate, average speed and average vehicle spacing
    of proposed model and a conventional vehicle formation model.
    Baseline DRL Model DRL + Formation Maintenance Proposed Model
    Collision rate(%) 61.25 5.03 5.40 3.00
    Average Speed(km/h) 59.09 73.87 74.08 74.98
    Average spacing(m) 5.92 5.17 3.10 3.01
  2. 2.
    Scenario Evaluation: We focused on traffic flow oscillations, a common and representative scenario, to assess the performance of the traditional CACC and CLCC, a reinforcement learning-based model, and our proposed model. We analyzed the time-varying behavior of headway, acceleration, speed, and lateral position under these conditions, as illustrated in Figure 9.
    Figure 9: Schematic illustration of headway, acceleration, velocity and lateral position of vehicle formations in three typical risk cases.
    Safety: The traditional CACC and CLCC, influenced by the IDM-lead vehicle, displayed significant sensitivity to the motion of preceding vehicles, leading to frequent adjustments in acceleration and speed. This resulted in unstable behavior, particularly during traffic flow oscillations. In contrast, both the reinforcement learning-based model and our proposed model executed adaptive lane changes by the 7th time step, effectively mitigating the impact of oscillations and enhancing overall safety.
    Efficiency: The traditional CACC and CLCC, lacking platoon reconfiguration capabilities, achieved efficiency gains only in the presence of large gaps. Concurrently, the intricate traffic patterns have precluded CLCC from identifying viable locations for collaborative lane changes, thereby impeding the potential for enhanced operational efficiency. Despite maintaining smaller headways, its lower speed reduced overall efficiency. The reinforcement learning-based model improved platoon speed by reconfiguring to more optimal lane positions. Our proposed model further enhanced efficiency by maintaining the platoon post-reconfiguration, sustaining high speeds and smaller headways, leading to more stable and efficient operation.
    To generalize our findings, we tested 30,000 randomly generated high-interference scenarios to evaluate three models: the DRL model, the DRL model with formation-keeping, and our complete proposed model, including the twin-world model. The baseline platoon consisted of an IDM lead vehicle with LQR-following vehicles maintaining a 0.5-second headway.
    In contrast, traditional CACC and CLCC models exhibit high collision rates in high-interference scenarios. However, our proposed method significantly reduces these rates while also enhancing efficiency. It is important to note that due to the randomness and complexity of generated scenarios, some unavoidable situations may still lead to collisions across all algorithms. As shown in Table II, our proposed model outperforms traditional strategies across multiple metrics. The DRL model effectively handles emergency scenarios, reducing the collision rate to below 5%. Incorporating formation-keeping maintains a comparable collision rate of 5.40%. However, the twin-world safety model further reduces the collision rate to 3%, underscoring the substantial impact of safety simulations.
    Moreover, the proposed model achieves an average platoon passing speed exceeding 70 km/h, a significant improvement over the baseline’s 59.09 km/h, indicating enhanced safety and optimized traffic flow. Additionally, the model supports smaller headways, achieving an average spacing of 3 meters compared to 5.92 meters with traditional methods, reflecting improved vehicle control and coordination.
    Overall, the proposed model outperforms traditional platooning strategies, demonstrating substantial improvements in safety, robustness, and efficiency.

### v2

@@ III-B Reward for Platooning
where $r_{C}$ penalizes collisions, $r_{L}$ encourages lane discipline, $r_{F}$ rewards maintaining high speeds, and $r_{A}$ controls acceleration behavior. The weights $w_{C}$, $w_{L}$, $w_{F}$, and $w_{A}$ are tuned to reflect the importance of each factor, with a strong emphasis on collision avoidance through a higher value of $w_{C}$.
From the platoon-wide perspective, the reward function promotes the cooperation and cohesion of the entire platoon. The system-wide reward $R_{sys}(s,a)$ encourages vehicles to remain coordinated and maintain proper formation:
@@
In this expression, $r_{M}$ rewards staying in the same lane, $r_{D}$ encourages covering greater distances, $r_{H}$ ensures proper headway between vehicles, and $r_{S}$ aligns speed coordination within the platoon. The corresponding weights $w_{M}$, $w_{D}$, $w_{H}$, and $w_{S}$ are designed to ensure that each vehicle’s actions contribute to the collective success of the formation.
This reward design effectively balances the individual objectives of safety and efficiency with the collective need for coordination, ensuring that the platoon operates as a cohesive unit. The combined rewards guide both individual vehicles and the platoon as a whole toward achieving optimal performance in dynamic traffic conditions. Figure 4 illustrates the structure of the reward function, highlighting how it integrates both individual and collective factors.
Figure 4: Illustration of the designed headway maintaining reward, which considers both the single-vehicle rewards and the multi-vehicle rewards.
III-C Deep Reinforcement Learning Algorithm
Traditional DRL algorithms, such as Deep Q-Networks (DQN), have been applied in autonomous driving decision-making tasks. However, these algorithms primarily focus on single-agent scenarios and often struggle in multi-agent cooperative decision-making environments, such as vehicle platooning. This study addresses the challenge of reconstructing vehicle platoons in mixed traffic, aiming to develop a robust self-organizing decision model capable of handling highly dynamic and disruptive conditions through collaborative agent behavior.
To address this challenge, TRPO is selected for its ability to maintain stability by constraining policy updates within a trust region, thus preventing instability caused by drastic changes. This characteristic makes it well-suited for cooperative multi-vehicle decision-making, where stable and incremental improvements are critical.
A CTDE strategy is employed, as shown in Figure 5, where sensor data from all platooning vehicles are aggregated into a unified observation space. A shared policy $\pi_{\theta}$ then generates actions for all vehicles within the platoon. Following the execution of these actions by the low-level controllers, the policy $\pi_{\theta}$ is updated based on environmental feedback using experience replay and gradient-based optimization.
Figure 5: Illustration of the TRPO-based formation configuration algorithm schematic model.
@@ III-C Deep Reinforcement Learning Algorithm
Deep neural networks (DNNs) are employed to approximate the policy and value functions, enabling agents to learn how to make decisions within our environment to maximize cumulative rewards. This study employs the Actor-Critic architecture-based DRL neural network structure depicted in Figure 6.
Figure 6: Illustration of the Actor-Critic Network Structure with Security-Enhanced Action Probability, which processes inputs through full conneted layers in the Actor-Critic networks. The Actor generates action probabilities, which are refined by a security mask to ensure safe decisions, while the Critic evaluates actions to maximize cumulative rewards.
@@
III-D Twin World Security Projection
While DRL enables robust decision-making in mixed traffic with strong interferences, its ”black-box” nature can lead to unexpected outcomes in standard scenarios and may not fully comply with traffic regulations. Although more reward functions could enforce these constraints, they may incur significant training costs and risk model divergence. These issues are addressed in this study through the implementation of a comprehensive and extensible safety mask.
To ensure decision safety and sample efficiency, a safety-priority-based twin-world model is proposed. This model evaluates the safety of DRL decisions by first detecting potential collisions based on safety priorities, thereby reducing computational complexity. It then recursively computes $T_{n}$ steps in a twin world constructed from the current scenario, optimizing actions based on the platoon’s safety margin.
Given the focus of this study on decision-making in mixed traffic, predicting the driving behavior of HDVs is essential. The prevalent Intelligent Driver Model (IDM) [29] is utilized for simulating car-following behaviors, while the MOBIL model [30] is adopted for lane-changing decisions. The combination of these models minimizes collision risk among HDVs while closely mirroring human driving patterns.
The safety mask models each vehicle as a circle, with the center at the vehicle’s center and a radius equal to half its length. Future $T_{n}$ steps are simulated using motion models of HDVs and CAVs to assess collision risks. While this simulation is straightforward for individual vehicles, it becomes computationally challenging for platoons due to the exponentially growing action space. To address this, a safety-priority-based simulation model is proposed, which prioritizes vehicles with the smallest safety margins and conducts safety simulations in the twin world, ensuring efficient real-time computation and platoon cohesion.
@@ 23      Replace action $a_{t}\leftarrow a_{t}^{\prime}$ Update $\phi[i]\leftarrow\phi[i+1],i=1,2,\cdots$
III-E Formation maintenance
Input: $vehicles$, $riskLevel$, $Participants$, $L_{safe}$
Output: $DecisionStrategy$
1
2for $i=0$ to $N$ do
3       Assess risk level and number of traffic participants within $L_{safe}$;
5             Switch to State $S_{1}$: LQR Control Model;
6             Activate LQR Control;
7             Set $DecisionStrategy=LQR$;
8
9      else
10             Switch to State $S_{2}$: Safety-Enhanced Data-Driven Strategy;
11             Set $DecisionStrategy=DataDriven$;
12
13       end if
14      Execute $DecisionStrategy$ based on the current state;
15       if $DecisionStrategy==LQR$ then
16             DecideWithLQR();
17
18       end if
19      else if $DecisionStrategy==DataDriven$ then
20             DecideWithDataDrivenStrategy();
21
22       end if
23
24 end for
return $DecisionStrategy$
Algorithm 2 Adaptive Switching Mechanism
The platoon reconfiguration model ensures safe, robust decision-making in complex, high-interference scenarios but struggles with maintaining close headway distances in routine conditions due to conservative safety strategies. To address this, an adaptive switching mechanism, the platoon maintenance model, is proposed, incorporating optimal control methods.
This model maintains close headway in routine scenarios while dynamically adjusting strategies based on real-time traffic conditions and interference levels, enhancing both safety in complex scenarios and efficiency in routine conditions.
To allocate decision strategies across scenarios, a Finite State Machine (FSM) $M(Q,D,\delta,q_{0},F)$ is used, where:
@@ 20             DecideWithDataDrivenStrategy();
This study considers a scenario as safe when the platoon is on the same lane and there are no other traffic participants within $L_{safe}$ ahead or behind the platoon. In this case, the FSM switches to state $S_{1}$, adopting the LQR control model to ensure smaller headway distances. In this study, string stability within the platoon[6] is not emphasized, as it has been thoroughly addressed in previous research. In all other scenarios, the FSM switches to state $S_{2}$, which employs a safety-enhanced data-driven strategy to balance the safety and efficiency of the platoon. By defining these states and transitions, the FSM provides a structured approach to adaptively manage the decision strategies of the platoon, ensuring optimal performance across varying traffic conditions. The detailed procedure is illustrated in Algorithm 2.
@@ IV-A1 Environment and simulation setting
The simulation environment is constructed upon the Highway-Env simulator[31], a reinforcement learning experimental environment library based on an OpenAI Gym environment. As illustrated in Figure 2, all vehicles’ high-level policies are transformed into throttle and steering information through a low-level PID controller, thereby enabling vehicle movement within the simulation environment. Furthermore, our HDVs employ the IDM[29] car-following model and the MOBIL[30] lane-changing model provided by Highway-Env to achieve effective control in both longitudinal and lateral directions.
During the training process, a varying number of HDVs will randomly appear at randomly generated refresh points. The sampling frequency is set to $15$ Hz, meaning decisions and controls are made 15 times per second.
Three different random seeds are used to train the proposed model in various complex scenarios, with each session running for over 100,000 steps until convergence. Four algorithms—DQN [32], A2C [33], PPO [27], and RecurrentPPO [34]—are selected as experimental baselines. The training parameters for the DRL model are detailed in Table I.
@@ IV-A2 Performance evaluation
This section evaluates the performance of the proposed model within a simulated training environment, benchmarking it against advanced RL algorithms and conventional platooning strategies. Our findings indicate that the proposed platooning model consistently outperforms existing approaches across a variety of challenging scenarios.
@@
The scenarios, depicted in Figure 1, are designed as edge cases to test the model’s safety and robustness under extreme conditions. The platoon vehicles travel at a speed of 28 m/s within a mixed traffic flow, where surrounding vehicles’ speed distribution follows a uniform range of 20 to 28 m/s. The traffic flow is characterized by oscillations and diverse driving styles, including aggressive, conservative, and neutral behaviors, creating a volatile environment prone to accidents.Additionally, malfunctioning vehicles may stall or brake unexpectedly, further increasing risk. The model is trained in this complex, high-risk scenario, with randomly generated traffic flows simulating real-world conditions.
Comparison with DRL baselines. Figure 7 illustrates the training performance of our model compared to baseline DRL algorithms across various scenarios. The training curves demonstrate that our model consistently outperforms the baseline algorithms, showing high adaptability and robustness essential for autonomous driving. This stability is attributed to our safety deduction model and innovative network structure, which enable our model to maintain reliable learning performance under challenging conditions and converge to a higher reward level.
Table II provides specific performance data. While our model may occasionally exhibit slightly lower speed and greater average headway distance compared to some baselines, it achieves significantly higher rewards, indicating that it prioritizes safety over raw efficiency. In contrast, the baseline algorithms reach their performance metrics at the expense of safety, which has been further verified in simulation experiments. Such trade-offs are unacceptable in transportation systems where safety is paramount.
In scenarios with high randomness and interference, such as traffic flow oscillations and traffic accidents, our model achieves significantly better convergence than other algorithms. This indicates an enhanced ability to learn optimal decisions under strong, dynamic disturbances, underscoring the robustness of our approach.
TABLE II: Performance metrics comparison between proposed method and benchmarks
Scenario Model Reward Avg. Speed Avg. HWD
Sce-1 A2C 18.93 22.23 10.00
DQN 9.47 25.18 20.06
PPO 22.10 23.83 13.85
RPPO 17.43 17.57 9.73
Ours 23.07 20.57 9.93
Sce-2 A2C 6.04 20.34 10.50
DQN 2.31 27.72 9.81
PPO 15.05 20.75 14.37
RPPO 4.10 25.70 17.10
Ours 31.66 20.59 11.33
Sce-3 A2C 4.87 20.18 10.00
DQN 11.83 20.71 12.18
PPO 58.21 20.13 10.00
RPPO 34.19 20.15 9.46
Ours 61.07 20.93 8.95
Comparison with Leading Platooning Algorithms. To evaluate the effectiveness of the TriCoD framework, a comparative analysis is conducted with several leading platooning algorithms. This analysis is based on four key performance indicators: average vehicle speed, average headway, collision rate, and pass rate. These metrics provide a comprehensive assessment of the model’s performance in terms of safety and efficiency. Notably, the pass rate metric is designed to consider scenarios where the convoy may need to halt to avoid hazards, thereby balancing risk avoidance with maintaining traffic flow.
In the experiments, the model is evaluated against advanced lane-changing algorithms, including SuPLC[35] and SiPLC[36], as well as a baseline DRL model across different scenarios: Human Inference, Traffic Accidents, and Traffic Flow Oscillation. As shown in Table III, our model consistently outperformed these benchmarks in terms of collision rate and pass rate, demonstrating a unique balance between safety and operational efficiency.
While some baseline models achieved slightly higher speeds or tighter headway distances, they do so at the expense of safety, as indicated by increased collision rates. In contrast, the TriCoD model maintained a zero-collision rate and a perfect pass rate of 100%, while still being able to maintain a relatively high pass rate and a smaller average headway. This balance illustrates our model’s capacity to prioritize safety without sacrificing operational effectiveness, especially in scenarios involving high-risk factors like traffic accidents and flow oscillations.
Overall, these results indicate that our model consistently achieves high safety and operational metrics across varied and challenging environments, demonstrating an effective balance that baseline models do not maintain.
TABLE III: Performance metrics comparison of multiple metrics
between the proposed method and three state-of-the-art benchmarks
Scenario Model Avg. Speed[m/s] Avg. HWD[m] Collision Rate Pass Rate
Human Inference SuPLC 22.58 9.68 1.00 0.00
SiPLC 22.58 9.68 1.00 0.00
DRL 20.46 10.00 0.00 1.00
Ours 20.57 9.93 0.00 1.00
Traffic Accidents SuPLC 21.96 10.18 0.52 0.29
SiPLC 20.81 10.15 0.54 0.21
DRL 21.13 14.10 0.19 0.91
Ours 20.59 11.33 0.00 1.00
Traffic Flow Oscillation SuPLC 20.28 10.13 0.03 1.00
SiPLC 19.68 10.15 0.03 0.97
DRL 20.93 10.08 0.09 0.91
Ours 20.93 8.95 0.00 1.00